OpenStudy (anonymous):

A circle has a diameter of thirty-two yards. What is its area? Use three point one four for pi.

OpenStudy (anonymous):

@ArtemisWolf @joannacaro123 @Poop_Emoji

OpenStudy (anonymous):

I'll give a medal

OpenStudy (anonymous):

32 * 3.14 = ?. I know you don't want the answer.

OpenStudy (anonymous):

How to calculate the diameter of a circle: d=2r How to calculate circumference: c=2·π·3 How to calculate radius: r= C 2·π How to calculate area in a circle: π(rr)

OpenStudy (anonymous):

Or do you?

OpenStudy (anonymous):

Thanks you know I don't

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

u got a medal :)

OpenStudy (anonymous):

Thnx. So did u! :)